# Knowledge Circuits in Pretrained Transformers

### v1

@@ 5 Knowledge Circuits Elucidate Internal Mechanisms for Knowledge Editing
In this section, our objective is to evaluate the impact of previous knowledge editing methods (why fail in certain cases and settings) and validate the effectiveness of knowledge circuits.
@@
Here, we adopt the ROME method [18] and FT-M [24], which aim to edit the MLP layers in the language model. The most important hyper-parameter in knowledge editing is the layer, as the same method’s performance varies significantly via the layers. Here, we evaluate the performance of different editing layers and their effectiveness. We compare the knowledge circuits computed by the edited model with the original one, and we present results in Figure 4 and report details in Appendix D. As discussed in the previous part, the early-to-middle layers are the main part of aggregating the target entity $o$ to the top rank. In the original model, the probability of the target entity “Intel” is nearly zero, and the model fails to elevate it to the top rank in the vocabulary. Editing layer 00 with ROME and FT-M both give us the correct answer but we can view different scenarios for their knowledge circuits. For ROME, as the correct information is added to the subject position, we can recognize a behavior of the Mover Head shifts from copying to extracting the edited information from the subject position. This information gradually aggregates through the subsequent layers, and by layer 15, “Intel” emerges as the top-ranked entity, with its probability increasing significantly. Specially, before editing, the mover head L15H3 attends to the “controller” token and returns “controller” as the output, while in the edited model, the attention head’s output moves to the “Intel”, which means the model gains the information at the subject space. While for FT-M, the edited model tends to directly write the knowledge into the specific component, which would greatly dominate the following component in the model. As shown in Figure 4, the output logits in MLP-0 for “Intel” are more than 10, and it emerges as the top rank in the residual stream directly. This phenomenon can be found in different knowledge types and layers and we report results in Appendix D.2. However, the added knowledge may have the risk to influence unrelated knowledge. When we test another fact “Windows server”, the model still tends to give us the “Intel” answer, demonstrating the overfitting problem. This finding supports previous analysis regarding the correlation between localization and editing [45], suggesting that edits may not alter the storage but merely add signals into the knowledge circuits.
@@ 6 Knowledge Circuits Facilitate Interpreting Language Model Behaviors
We focus on factual hallucinations, which occur when the model provides an incorrect target entity for a given subject $s$ and relation $r$. In our experiments (Figure 5 and Appendix C.2), we observe that the model fails to move the correct knowledge to the final token in the earlier layers. This failure is evident as the circuit lacks an effective mover head or the mover head selects incorrect information. For instance, in the prompt “The official currency of Malaysia is called”, both the correct answer “Ringgit” and the incorrect one “Malaysian” are accumulated before layer 15. However, at layer 16, the mover head L15H10 extracts the erroneous information. Despite a rank drop of the true one in layers 20–22, this is insufficient to correct the previous mistake.
@@
Despite storing vast amounts of knowledge, a language model may still provide incorrect answers. However, with demonstrations or examples (based on RAG [49]), it can quickly generate correct responses. To this end, we focus on the scenario where the model initially provides an incorrect answer but can then produce the correct response upon receiving the appropriate demonstration. We consider the original knowledge circuit and introduce a new knowledge circuit based on the demonstration. Our analysis reveals that, compared to the zero-shot knowledge circuit, several new attention heads appear in the computation graph when the demonstration is incorporated as shown in Figure 5. These heads mainly focus on the demonstration’s context: “The co mparative form of small is smaller”. Concretely, Todd et al. [50] have identified a concept known as the Function Vector, which represents the average of some key attention heads. In our experiments, we found that these attention heads primarily focus on the demonstration, indicating their role in leveraging the demonstration to reactivate and correct the model’s response.
@@ 7 Related Work
Knowledge Mechanism of Transformers. How the language model store and utilize knowledge is an ongoing research topic. Previous works find that the MLP in Transformers works as a key-value memory and stores enormous knowledge [12, 15, 14, 18]. As to the relation between entities, Hernandez et al. [42] observe that facts can be decoded linearly from the enriched residual stream of the subject by mapping the subject entity to the object entity. Instead of viewing the knowledge storage in isolation, Geva et al. [13], Lv et al. [31], Yu and Ananiadou [16] find the knowledge is accumulated during the layers. Regarding knowledge analysis, Bayazit et al. [51] also attempt to discover critical knowledge in language models. However, they only consider several layers in the model and use the pruning method, which may overlook the connections between components. More related works can be found in Appendix E.1.
Manipulate Language Models. Recently, many works aim to manipulate the language models to make the model aligned with world knowledge or social value norms, such as knowledge editing [20, 24], machine unlearning [52, 53] and detoxification [54, 55]. Most of these works are elicited by previous knowledge mechanism findings such as knowledge neuron [56]. They modify the MLP in the LLM [18, 12] to change the model’s behavior based on specific factual knowledge. However, recent works [57, 58] demonstrate the pivotal role of the attention part in knowledge representation. Hase et al. [45] also observe that the performance of editing within a layer may not reliably pinpoint the location of the fact. In this paper, we try to manipulate specific knowledge of language model via knowledge circuit, including both MLP and attention components across different layers.
@@ 8 Conclusion
In this paper, we present a new perspective on knowledge storage based on circuit theory and conduct a preliminary analysis to demonstrate its effectiveness. We hope these findings can advance our understanding of the knowledge mechanisms of language models and provide insights for better designing and editing language models, enhancing knowledge, and improving reasoning to enhance factuality and alleviate hallucinations.
@@
Current circuit discovery-based patching method is time-consuming, there are some concurrent works that propose more efficient way [59] to build the model’s information flow. Also, there are some other methods to discover circuits, like acdcpp [60] and Sparse Auto-Encoders [61, 36]. We believe that knowledge circuit discovery has a huge room for improvement. Additionally, by focusing on linguistic, factual, commonsense, and bias-related knowledge, we believe our approach can be applied to ensure safety and privacy information to promote trustworthy AI.
@@ Appendix A Implementation Details
All the data used in our paper is sourced from Hernandez et al. [42], with detailed information provided in Table 5. In the original setting, they use the data for few-shot settings, but in our experiments, we consider zero-shot knowledge storage, so here we sample the data using the Hit@10 to detect whether the model understands knowledge for the given prompt based on the $s$ and $o$. We sample the test set in a 1:1 ratio with the validation set to ensure a balanced evaluation.
@@ B.2 Special Components in Knowledge Circuit
When zooming into the discovered circuit, we can find several kinds of special attention heads, or MLPs, in the model that play a pivotal role in the final prediction, similar to what previous research has indicated [31]. We follow previous work [31, 17] and list the definition of these heads as follows:
  • Mover Head [31, 27]: These heads focus on the last token of the context and attend to the subject token, functioning as a mover to transfer information.
  • Relation Head [17]: These heads would attend to the relation token in the context and produce some relation-related tokens that would guide the behaviour of the following components.
  • Mix Head [17, 43]: These head would focus on both the relation token and the subject tokens. In our experiments, we found these heads usually work similar as the mover head.
@@
In our experiment, we view the model as a collaboration between different nodes in the identified circuit. They perform as different kind of attention heads, like mover head and relation head. We list some heads that are responsible for the storage of different kind of knowledge and relations in Table 3.
@@
Merullo et al. [27] have identified shared circuits for the IOI task and the Colored Objects task. We also observe this phenomenon in the factual recall task. As depicted in Table 3, we can observe that for related relations such as “city_in_country”, “name_birth_place”, and “country_language”, their circuits include both L21H12, which stores and maps country-related information. Additionally, we found that some relation heads are activated by different relations. For instance, in our experiments, the head ‘L7H14’ appears in the circuits of both “official_language” and “official_currency”. We speculate that these reused heads, rather than task-specific heads, can be considered as topic heads, as proposed by [66, 59]. We believe that further investigation into this distinction is warranted in future research.
@@ B.3 A Case on TinyLLAMA
Figure 7: A simplified knowledge circuit found in TinyLLAMA for the knowledge “The mother tongue of Thierry Mugle is French”.
@@ C.1 Multi-hop Factual Knowledge Editing
Actually, there are two ways for the model to conduct multi-hop reasoning. As shown in the figure, the model can also answer the question by combining the two hop relations together (in the given case, combine “hometown” and “language” as “mother tongue”) If the model is capable of combining two-hop relations in a more integrated or semantic way, such as inferring that the “mother tongue” is the language of one’s hometown, this suggests a more complex reasoning process that goes beyond the simple overlap of nodes and edges. To capture this kind of reasoning, we assesses the model’s ability to integrate information from different hops. $R_{integrated}$ as the set of integrated relations (new paths created by combining information from different hops)
@@
It further enhances our previous findings that the model actually conducts the factual recall with the relation head and the gathered information about the subject. Moreover, we found this phenomena is hugely alleviated by the aligned model, which requires further investigation in the future.
@@ C.3 Reverse Relation
Reverse Curse [68] is an important issue in language models when models successfully give us the correct answer $o$ for $(s,r)$, but with a reverse relation $\hat{r}$ and $o$, the models fail to give us the correct subject $s$. In this part, we endeavor to investigate how the language model manages the reverse relation when they successfully store the knowledge. We first sample facts where the language model successfully predicts the given knowledge and the reversed fact. Then, we compute the overlap between these two circuits, $\mathcal{C}_{d}$ and $\mathcal{C}_{r}$ under node levels based on equation 4. We select the “superhero_person” relation to see the difference of these two circuit and test the node overlap of the two circuit in the model. We notice that the overlap between the two circuits is about 70%, indicating the language model may store the related information in the same place. We also found the activated mover heads for the two relationships to be identical.
@@ D.1 Method Implementation
As proposed by Meng et al. [18], ROME view knowledge editing as a minimal optimization problem. ROME regards the MLP module as a simple key-value store. Specifically, the key represents a subject and the value encapsulates knowledge about that subject, the MLP can reestablish the association by retrieving the corresponding value for the key. To add a new key-value pair, ROME applies a rank-one modification to the MLP’s weights, effectively “writing in” the new information directly. This method enables more direct and precise modification of the model’s knowledge. The ROME method for model editing was conducted based on the EasyEdit[69] framework, utilizing the default parameters provided by EasyEdit. The experiments are performed on an A800 80G GPU, with approximately 8GB of memory consumption.
@@
For Fine-Tuning (FT-M), we follow Zhang et al. [24]. It trains the MLP layer using the cross-entropy loss on the target answer while masking the original text. This approach aligns more closely with the traditional fine-tuning object. The FT-M method is conducted using the EasyEdit [^4] https://github.com/zjunlp/EasyEdit [69] framework, with the default parameters provided by EasyEdit. The experiments are also performed on an A800 80G GPU, with memory consumption of approximately 10GB.
@@ D.2 Edit Cases on FT-M and ROME
When a circuit is established for a particular piece of knowledge, we can manipulate the model’s computation by targeting critical points within the circuit. Li et al. [70] ablates a small number of important causal pathways by masking the edges in the circuit and making the model less toxic and safer, which proves the effectiveness of the circuit. As illustrated in figure 11 and figure 12, we present the changes in the ranking of the predicted probabilities for the target new token when editing layer 6, 12, and 18 of the GPT-2 medium model using FT-M and ROME methods. When applying FT-M for model editing, it is evident that the rank of the target new token’s probability sharply declines at the corresponding edited layer, resulting in a vertical line in the figure. This indicates that FT-M directly embeds the editing information into the model’s information flow. Conversely, when using the ROME method for editing, this effect is mitigated. The predicted probability of target new token reaches its peak only a few layer after the edited layer. This observation is consistent with our previous analysis in Section 5.

### v2

@@ 5 Knowledge Circuits Elucidate Internal Mechanisms for Knowledge Editing
In this section, our objective is to evaluate the impact of previous knowledge editing methods and validate the effectiveness of knowledge circuits. We aim to understand why these methods may fail in certain cases and settings, which can also help deepen the understanding of the knowledge circuit.
@@
Here, we adopt the ROME method [18] and FT-M [24], which aim to edit the MLP layers in the language model. The most important hyper-parameter in knowledge editing is the layer, as the same method’s performance varies significantly via the layers. Here, we evaluate the performance of different editing layers and their effectiveness. We compare the knowledge circuits computed by the edited model with the original one, and we present results in Figure 4 and report details in Appendix D. As discussed in the previous part, the early-to-middle layers are the main part of aggregating the target entity $o$ to the top rank. In the original model, the probability of the target entity “Intel” is nearly zero, and the model fails to elevate it to the top rank in the vocabulary. Editing the model with ROME and FT-M both give us the correct answer but we can view different scenarios for their knowledge circuits. For ROME, as the correct information is added to the subject position, we can recognize a behavior of the Mover Head shifts from copying to extracting the edited information from the subject position. This information gradually aggregates through the subsequent layers, and by layer 15, “Intel” emerges as the top-ranked entity with its probability increasing significantly. Specially, before editing, the mover head L15H3 attends to the “controller” token and returns “controller” as the output, while in the edited model, the attention head’s output moves to the “Intel”, which means the model gains the information at the subject space. For FT-M, the edited model tends to directly write the knowledge into the specific component, which would greatly dominate the following component in the model. As shown in Figure 4, the output logits in MLP-0 for “Intel” are more than 10, and it emerges as the top rank in the residual stream directly. This phenomenon can be found in different knowledge types and layers and we report results in Appendix D.2. However, the added knowledge may have the risk of influencing unrelated knowledge. When we test another fact “Windows server”, the model still tends to give us the “Intel” answer, demonstrating the overfitting problem. This finding supports previous analysis regarding the correlation between localization and editing [45], suggesting that edits may not alter the storage but merely add signals into the knowledge circuits.
@@ 6 Knowledge Circuits Facilitate Interpreting Language Model Behaviors
If the knowledge is stored and expressed by the circuit we discovered, we aim to discover what happened when the model gave us the incorrect answer. We focus on factual hallucinations, which occur when the model provides an incorrect target entity for a given subject $s$ and relation $r$. In our experiments (Figure 5 and Appendix C.2), we observe that the model fails to move the correct knowledge to the final token in the earlier layers. This failure is evident as the circuit lacks an effective mover head or the mover head selects incorrect information. For instance, in the prompt “The official currency of Malaysia is called”, both the correct answer “Ringgit” and the incorrect one “Malaysian” are accumulated before layer 15. However, at layer 16, the mover head L15H10 extracts the erroneous information. Despite a rank drop of the true one in layers 20–22, this is insufficient to correct the previous mistake.
@@
Despite storing a vast amount of knowledge, a language model may still provide incorrect answers. However, with demonstrations or examples (based on RAG [49]), it can quickly generate correct responses. To this end, we focus on the scenario where the model initially provides an incorrect answer but can then produce the correct response upon receiving the appropriate demonstration. We consider the original knowledge circuit and introduce a new knowledge circuit based on the demonstration. Our analysis reveals that, compared to the zero-shot knowledge circuit, several new attention heads appear in the computation graph when the demonstration is incorporated. We show the behavior of these attention heads in Figure 5. We can see these heads mainly focus on the demonstration’s context: “The comparative form of small is smaller” and works as the Induction Head [50] that look back over the sequence for previous instances of the current token and find the token that came after it last time. To better view the function of these heads, we conduct experiments by ablating the newly appeared attention head in the ICL circuit in Table 2. We find that compared to the randomly selected attention head by ablating this attention, the probability drops significantly in the prediction, demonstrating the importance of these identified attention heads. These aligned with previous work where Todd et al. [51] have identified a concept known as the Function Vector, which represents the average of some key attention heads and provides the task learning ability.
Table 2: Performance change via ablating the newly appeared attention heads in the ICL circuit and random heads.
Knowledge Origin Model Ablating extra head Ablating random head
Linguistic adj_comparative 62.24 32.55 58.18
Commonsense word_sentiment 89.02 55.50 88.61
substance_phase 78.74 52.85 71.24
Bias occupation_gender 86.97 59.54 86.54
Factual person_occupation 35.17 23.27 31.60
@@ 7 Related Work
Knowledge Mechanism of Transformers. How the language model stores and utilizes knowledge is an ongoing research topic. Previous works find that the MLP in Transformers works as a key-value memory and stores enormous knowledge [12, 15, 14, 18]. As to the relation between entities, Hernandez et al. [42] observes that facts can be decoded linearly from the enriched residual stream of the subject by mapping the subject entity to the object entity. Instead of viewing the knowledge storage in isolation, Geva et al. [13], Lv et al. [31], Yu and Ananiadou [16] find the knowledge is accumulated during the layers. Regarding knowledge analysis, Bayazit et al. [52] also attempts to discover critical knowledge in language models. However, they only consider several layers in the model and use the pruning method, which may overlook the connections between components. More related works can be found in Appendix E.1.
Manipulate Language Models. Recently, many works aim to manipulate the language models to make the model aligned with world knowledge or social value norms, such as knowledge editing [20, 24], machine unlearning [53, 54] and detoxification [55, 56]. Most of these works are elicited by previous knowledge mechanism findings such as knowledge neuron [57]. They modify the MLP in the LLM [18, 12] to change the model’s behavior based on specific factual knowledge. However, recent works [58, 59] demonstrate the pivotal role of the attention part in knowledge representation. Hase et al. [45] also observe that the performance of editing within a layer may not reliably pinpoint the location of the fact. In this paper, we try to manipulate specific knowledge of language models via knowledge circuits, including both MLP and attention components across different layers.
@@ 8 Conclusion
In this paper, we present a new perspective on knowledge storage based on circuit theory and conduct a preliminary analysis to demonstrate its effectiveness. We found that knowledge circuits in the model are not only responsible for expressing knowledge but can also guide behavior in different settings. We hope these findings can advance our understanding of the knowledge mechanisms of language models and provide insights for better designing and editing language models, enhancing knowledge, and improving reasoning to enhance factuality and alleviate hallucinations. In the future, we may further explore how these attention heads are formed during pretraining and activated.
@@
In this work, we employ the causal mediation method to automatically construct circuits tailored to specific knowledge domains. However, this circuit discovery-based patching approach is time-intensive. Contemporary research efforts have introduced more efficient methodologies for modeling information flow [60, 61, 62]. Additionally, alternative techniques for discovering circuits through mask training [63, 64] and Sparse Auto-Encoders [65, 36] have been proposed, highlighting diverse facets of circuit behavior within large language models (LLMs). We posit that the field of knowledge circuit discovery holds significant potential for advancement. Furthermore, recent studies [66] have developed ’circuit breakers’ to manage representations associated with potentially harmful outputs. By concentrating on linguistic, factual, commonsense, and bias-related knowledge, we assert that our methodology can contribute to ensuring the safety and privacy of information, thereby fostering the development of trustworthy AI.
Acknowledgements
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), SMP-Zhipu.AI Large Model Cross-Disciplinary Fund, Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ Appendix A Implementation Details
All the data used in our paper is sourced from Hernandez et al. [42], with the detailed information provided in Table 6. In the original setting, they use the data for few-shot settings, but in our experiments, we consider zero-shot knowledge storage, so here we sample the data using the Hit@10 to detect whether the model understands knowledge for the given prompt based on the $s$ and $o$. We sample the test set in a 1:1 ratio with the validation set to ensure a balanced evaluation.
@@ B.2 Special Components in Knowledge Circuit
When zooming into the discovered circuit, we can find several kinds of special attention heads, or MLPs, in the model that play a pivotal role in the final prediction, similar to what previous research has indicated [31]. Apart from mover head and relation head, there is another kind of head named Mix Head [17, 43] which would focus on both the relation token and the subject tokens. In our experiments, we found these heads usually work similarly to the mover head. In particular, mover heads contribute more to the subject-specific information, as ablating them would increase other relation-related tokens’ probability. Moreover, if we ablate the relation head, we can find the model tends to generate some meaningless tokens instead of the relation information. From Figure 3 6, we can find ablating the mover head would increase the probability of "Italian", "English" and "Spanish", which are not subject-related. While ablating the relation head would lead to the increase of some meaningless words "a", and "that", which are not relation-related.
Figure 6: The output of the model. Ablating the mover head would increase the probability of "Italian", "English" and "Spanish", which are not subject-related. While ablating the relation head would lead to the increase of some meaningless words "a", "that", which are not relation-related.
@@
In our experiment, we view the model as a collaboration between different nodes in the identified circuit. They perform as different kinds of attention heads, like mover head and relation head. We list some heads that are responsible for the storage of different kinds of knowledge and relations in Table 4.
@@
Merullo et al. [27] have identified shared circuits for the IOI task and the Colored Objects task. We also observe this phenomenon in the factual recall task. As depicted in Table 4, we can observe that for related relations such as “city_in_country”, “name_birth_place”, and “country_language”, their circuits include both L21H12, which stores and maps country-related information. Additionally, we found that some relation heads are activated by different relations. For instance, in our experiments, the head ‘L7H14’ appears in the circuits of both “official_language” and “official_currency”. We speculate that these reused heads, rather than task-specific heads, can be considered topic heads, as proposed by [71, 60]. We believe that further investigation into this distinction is warranted in future research.
@@ B.3 A Case on TinyLLAMA
Figure 8: A simplified knowledge circuit found in TinyLLAMA for the knowledge “The mother tongue of Thierry Mugler is French”.
@@ C.1 Multi-hop Factual Knowledge Editing
Actually, there are two ways for the model to conduct multi-hop reasoning. As shown in the figure, the model can also answer the question by combining the two hop relations together (in the given case, combine “hometown” and “language” as “mother tongue”) If the model is capable of combining two-hop relations in a more integrated or semantic way, such as inferring that the “mother tongue” is the language of one’s hometown, this suggests a more complex reasoning process that goes beyond the simple overlap of nodes and edges. To capture this kind of reasoning, we assess the model’s ability to integrate information from different hops. $R_{integrated}$ as the set of integrated relations (new paths created by combining information from different hops)
@@
It further enhances our previous findings that the model actually conducts the factual recall with the relation head and the gathered information about the subject. Moreover, we found this phenomenon is hugely alleviated by the aligned model, which requires further investigation in the future.
@@ C.3 Reverse Relation
Reverse Curse [73] is an important issue in language models when models successfully give us the correct answer $o$ for $(s,r)$, but with a reverse relation $\hat{r}$ and $o$, the models fail to give us the correct subject $s$. In this part, we endeavor to investigate how the language model manages the reverse relation when they successfully store the knowledge. We first sample facts where the language model successfully predicts the given knowledge and the reversed fact. Then, we compute the overlap between these two circuits, $\mathcal{C}_{d}$ and $\mathcal{C}_{r}$ under node levels based on equation 4. We select the “superhero_person” relation to see the difference between these two circuits and test the node overlap of the two circuits in the model. We notice that the overlap between the two circuits is about 70%, indicating the language model may store the related information in the same place. We also found the activated mover heads for the two relationships to be identical.
@@ D.1 Method Implementation
As proposed by Meng et al. [18], ROME views knowledge editing as a minimal optimization problem. ROME regards the MLP module as a simple key-value store. Specifically, the key represents a subject and the value encapsulates knowledge about that subject, the MLP can reestablish the association by retrieving the corresponding value for the key. To add a new key-value pair, ROME applies a rank-one modification to the MLP’s weights, effectively “writing in” the new information directly. This method enables more direct and precise modification of the model’s knowledge. The ROME method for model editing was conducted based on the EasyEdit[74] framework, utilizing the default parameters provided by EasyEdit. The experiments are performed on an A800 80G GPU, with approximately 8GB of memory consumption.
@@
For Fine-Tuning (FT-M), we follow Zhang et al. [24]. It trains the MLP layer using the cross-entropy loss on the target answer while masking the original text. This approach aligns more closely with the traditional fine-tuning object. The FT-M method is conducted using the EasyEdit [^4] https://github.com/zjunlp/EasyEdit [74] framework, with the default parameters provided by EasyEdit. The experiments are also performed on an A800 80G GPU, with a memory consumption of approximately 10GB.
@@ D.2 Edit Cases on FT-M and ROME
When a circuit is established for a particular piece of knowledge, we can manipulate the model’s computation by targeting critical points within the circuit. Li et al. [75] ablates a small number of important causal pathways by masking the edges in the circuit and making the model less toxic and safer, which proves the effectiveness of the circuit. As illustrated in figure 12 and figure 13, we present the changes in the ranking of the predicted probabilities for the target new token when editing layer 6, 12, and 18 of the GPT-2 medium model using FT-M and ROME methods. When applying FT-M for model editing, it is evident that the rank of the target new token’s probability sharply declines at the corresponding edited layer, resulting in a vertical line in the figure. This indicates that FT-M directly embeds the editing information into the model’s information flow. Conversely, when using the ROME method for editing, this effect is mitigated. The predicted probability of the target new token reaches its peak only a few layers after the edited layer. This observation is consistent with our previous analysis in Section 5.